OpenStudy (anonymous):

What is the slope of the line which passes through (4, 5) and (0, 1)? A. 4 B. 1 C. undefined D. 0

OpenStudy (anonymous):

b.1

OpenStudy (karatechopper):

y-y/x-x

OpenStudy (anonymous):

Thank you<33

OpenStudy (amistre64):

finding the slope between 2 points is as simple as subtracting the points and stacking y/x

OpenStudy (amistre64):

(4, 5) -(0, 1) ------ 4, 4 ; y/x = 4/4 = 1